to review a method by which this problem can be solved. The electric field has a constant value of 2.8 x 102 V/m and is directed downward. The fieid is the same everywhere. The potential at a point P within this region is 121 V. Find the potential at the following points: (0) 9.5 x 103 m directy above P, (b) 3.9 x 103 m directly below P, (6) 5.3x 103m directly to the right of P (a) Number (e) Number By accessing this Question Assistance, you will learn while you earn points based on the Point Potential Policy set by your instructor Question Attempts: 0 of 3 used SAVE FOR LATE SUBHIT ANSWER
Answer #1

a) V' = V + E d = 121 + (280 * 9.5 * 10-3)

= 123.66 V

b) V'' = V - E d' = 121 - (280 * 3.9 * 10-3)

= 119.9 V

c) It has same potential as P , V''' = V

= 121 V